ABC is a right angled isosceles triangle right angled at A (2,1) . If the equation of side BC is 2x+y = 3 then the combined equation of line AB and AC is

Grade:

1 Answers

Arun
5 years ago

slope of BC= -2

slope of AB=3
slope of AD=1/2
slope of AC=-1/3

eqn of AB is y-3x+5=0
eqn of AD is 2y-x=0
eqn of AC is 3y+x-5=0

Multiply, two lines at a time for the three answers